OpenStudy (anonymous):

ind the derivative of f(x) = -4x2 + 11x at x = 10.

OpenStudy (anonymous):

i got -300 am i right?

OpenStudy (danjs):

one rule for the powers \[\frac{ d }{ dx }[a*x^n] = a*(n)x^{n-1}\]

OpenStudy (danjs):

The derivative of f(x) , at the point (10 , f(10)) , is the slope of a tangent line of f(x) through that point.

OpenStudy (anonymous):

wait so am i wrong?

OpenStudy (danjs):

what did you get for the derivative f ' (x)

OpenStudy (anonymous):

well i got -300 as my final answer

OpenStudy (danjs):

can you show the calculation to get that ...

OpenStudy (anonymous):

i honestly dont remeber, i think i used the (x+h)/h formula thing